A girl with a mass of 45 kg is standing on a scale in a lift. What is the weight of the girl when (i) the lift is accelerating upwards at 2.0 ms s −2 and (ii) the lift is moving down with a constant velocity. Calculate the inertial force in each case.
1
Expert's answer
2019-05-15T05:47:50-0400

The weight of the girl in a lift

"W=m(g \pm a)"

(a)

Weight

"W=45\:\rm{kg}\times(9.8\:\rm{m/s^2}+2.0\:\rm{m/s^2})=531\:\rm{N}"


Inertial force


"F_i=W-mg=531\:\rm{N}-441\:\rm{N}=90\:\rm{N}"

(b)

Weight

"W=45\:\rm{kg}(9.8\:\rm{m/s^2}+0\:\rm{m/s^2})=441\:\rm{N}"

Inertial force


"F_i=W-mg=441\:\rm{N}-441\:\rm{N}=0\:\rm{N}"